# Improved calculation of radiative corrections to $\boldsymbol{\tau\to\pi\pi\nu_{\tau}}$ decays

### v1

@@ Abstract
and would allow for a test of the Standard Model at an unprecedented level of precision. Unfortunately, the limiting factor is represented by the precision of the current theory prediction Aliberti et al. (2025a); Aoyama et al. (2012); Volkov (2019, 2024); Aoyama et al. (2025); Parker et al. (2018); Morel et al. (2020); Fan et al. (2023); Czarnecki et al. (2003); Gnendiger et al. (2013); Lüdtke et al. (2025); Hoferichter et al. (2025a); Blum et al. (2018); Giusti et al. (2019); Borsányi et al. (2021); Lehner and Meyer (2020); Wang et al. (2023); Aubin et al. (2022); Cè et al. (2022); Alexandrou et al. (2023); Blum et al. (2023); Kuberski et al. (2024); Boccaletti et al. (2024); Spiegel and Lehner (2025); Blum et al. (2025a); Djukanovic et al. (2025); Alexandrou et al. (2025); Bazavov et al. (2025a); Bazavov et al. (2025b); Keshavarzi et al. (2020); Di Luzio et al. (2025); Kurz et al. (2014); Colangelo et al. (2015); Masjuan and Sánchez-Puertas (2017); Colangelo et al. (2017a, b); Hoferichter et al. (2018a, b); Eichmann et al. (2019); Bijnens et al. (2019); Leutgeb and Rebhan (2020); Cappiello et al. (2020); Masjuan et al. (2022); Bijnens et al. (2020, 2021); Danilkin et al. (2021); Stamen et al. (2022); Leutgeb et al. (2023); Hoferichter et al. (2023a); Hoferichter et al. (2024); Estrada et al. (2024); Deineka et al. (2025); Eichmann et al. (2025); Bijnens et al. (2025); Hoferichter et al. (2025b, c); Holz et al. (2025a, b); Cappiello et al. (2025); Colangelo et al. (2014); Blum et al. (2020); Chao et al. (2021, 2022); Blum et al. (2025b); Fodor et al. (2025)
@@
which is worse than the experimental one by about a factor of four. The main source of theory uncertainty is the leading-order hadronic-vacuum-polarization contribution $a_{\mu}^{\text{HVP, LO}}$, which in Ref. Aliberti et al. (2025a) was estimated on the basis of lattice-QCD calculations, differing substantially from the previous $e^{+}e^{-}$-based result Aoyama et al. (2020); Davier et al. (2017); Keshavarzi et al. (2018); Colangelo et al. (2019); Hoferichter et al. (2019); Davier et al. (2020); Keshavarzi et al. (2020); Hoid et al. (2020). In fact, a data-driven estimate of this contribution was not even provided in Ref. Aliberti et al. (2025a) due to tensions in the $e^{+}e^{-}\to\text{hadrons}$ data base. The main rationale for this change was that after the CMD-3 measurement of the critical $e^{+}e^{-}\to\pi^{+}\pi^{-}$ channel Ignatov et al. (2024a, b) systematic tensions had increased to a level that could no longer be taken into account by a meaningful error inflation. Fortunately, there are a number of ongoing developments to rectify this situation Aliberti et al. (2025a); Colangelo et al. (2022a), including new data (such as the preliminary measurement by BaBar Pinto (2025)) as well as improvements of radiative corrections and Monte-Carlo generators Campanario et al. (2019); Ignatov and Lee (2022); Colangelo et al. (2022b); Monnard (2021); Abbiendi et al. (2022); Lees et al. (2023); Budassi et al. (2025); Aliberti et al. (2025b).
@@
A complementary avenue to shed light on the $2\pi$ contribution proceeds via $\tau\to\pi\pi\nu_{\tau}$ decays, in which case control over radiative corrections is arguably even more crucial. The $\tau$-based approach, pioneered in Ref. Alemany et al. (1998), relies on an isospin rotation from the $\pi^{\pm}\pi^{0}$ channel probed in $\tau$ decays back to $\pi^{+}\pi^{-}$ as required for the two-pion contribution $a_{\mu}^{\text{HVP, LO}}[\pi\pi]$. To make this work, first, the $\tau$-specific radiative corrections need to be removed from the measured spectral function and branching fraction, then isospin-breaking (IB) corrections need to be applied to the hadronic matrix elements, to account for the difference between the pion form factors $F_{\pi}^{V}(s)$ and $f_{+}(s)$ in the neutral and charged channel, respectively, and finally the $e^{+}e^{-}$-specific IB corrections need to be added back to match the conventional definition of $a_{\mu}^{\text{HVP, LO}}[\pi\pi]$. The latter can be extracted in a reliable way using a dispersive approach Colangelo et al. (2019, 2021); Colangelo et al. (2022c); Hoferichter et al. (2023b); Stoffer et al. (2023); Leplumey and Stoffer (2025) from the $e^{+}e^{-}\to\pi^{+}\pi^{-}$ data sets Achasov et al. (2006); Akhmetshin et al. (2007); Lees et al. (2012); Ablikim et al. (2016); Anastasi et al. (2018); Ignatov et al. (2024a), while the remaining IB corrections at present require model-dependent input Aliberti et al. (2025a), preventing the use of $\tau$ decays in the evaluation of $a_{\mu}^{\text{HVP, LO}}[\pi\pi]$.
@@
In this Letter, we address an improved calculation of the radiative corrections to $\tau\to\pi\pi\nu_{\tau}$, which presents a key challenge in the above program. So far, the evaluation of radiative corrections is based on chiral perturbation theory (ChPT) Cirigliano et al. (2001, 2002) in combination with resonance models Flores-Baéz et al. (2006, 2007); Davier et al. (2010); Miranda and Roig (2020); Davier et al. (2024); Castro et al. (2025), which makes robust uncertainty quantification difficult. Most importantly, we consider structure-dependent virtual corrections based on a dispersive representation of the pion form factor, which were previously shown to be sizable in the case of $e^{+}e^{-}\to\pi^{+}\pi^{-}$ Ignatov and Lee (2022); Colangelo et al. (2022b). In this context, we pay particular attention to the matching to ChPT, to be able to make the connection to short-distance (SD) contributions as well as lattice QCD Bruno et al. (2018). We also revisit the calculation of real corrections, including the singular threshold behavior, and develop a method for a stable numerical evaluation of these corrections down to the two-pion threshold. Finally, we perform fits to the available $\tau$ spectral functions from Belle Fujikawa et al. (2008), ALEPH Schael et al. (2005); Davier et al. (2014), CLEO Anderson et al. (2000), and OPAL Ackerstaff et al. (1999), to assess the consequences for the IB corrections in a $\tau$-based evaluation of the $2\pi$ channel, $a_{\mu}^{\text{HVP, LO}}[\pi\pi,\tau]$.
@@
Figure 1: Virtual photonic corrections to $\tau\to\pi\pi\nu_{\tau}$ in ChPT. Solid lines denote leptons, dashed lines pions, and wiggly lines photons. Not shown are self-energy diagrams and contact terms.
The SD contributions are subsumed into $S_{\text{EW}}^{\pi\pi}=1+2\alpha/\pi\log(M_{Z}/m_{\tau})+\cdots=1.0233(3)(24)$ Sirlin (1982); Marciano and Sirlin (1986, 1988, 1993); Braaten and Li (1990); Erler (2004); Davier et al. (2003); Cirigliano et al. (2023), where the uncertainty reflects the $\mathcal{O}(\alpha/\pi)$ scheme ambiguity pointed out in Ref. Aliberti et al. (2025a). $f_{+}(s)$ denotes the pion form factor in the charged current, normalized as $f_{+}(0)=1$, and radiative corrections are expressed in terms of the factor
@@
which is the central object of this study. It is normalized in such a way that $G_{\text{EM}}(s)=1+\mathcal{O}(\alpha)$, and amounts to averaging the amplitudes $f_{\text{loop}}(s,t)$, $g_{\text{rad}}(s,t)$ for virtual and real contributions, respectively, over the entire phase space. The variable $t$ represents the invariant mass squared of the $\tau$ and the charged pion, with integration boundaries $t_{\text{min}/\text{max}}(s)$, and the weight function reads
@@
This limitation can be overcome in a dispersive approach, in which case the diagrams in Fig. 1 need to be reinterpreted as unitarity diagrams. From this perspective, since diagrams $(b)$ and $(c)$ do not generate analytic singularities in the Mandelstam variables, only diagram $(a)$ needs to be considered further. Keeping in mind that the $\tau\nu_{\tau}\pi\pi^{0}$ vertex arises from integrating out the $W$ boson, this topology then takes the form shown in Fig. 2, and, accordingly, is referred to as “box diagram.” In general, its calculation requires knowledge of the full amplitude $\langle\pi\pi^{0}|j_{\text{em}}^{\mu}j_{W}^{\nu}|0\rangle$ of weak and electromagnetic currents, which could be constructed dispersively in analogy to pion Compton scattering García-Martín and Moussallam (2010); Hoferichter et al. (2011); Moussallam (2013); Danilkin and Vanderhaeghen (2019); Hoferichter and Stoffer (2019); Danilkin et al. (2020). In practice, by far the most dominant correction will arise from the pion-pole contribution indicated in Fig. 2. Following the strategy of Ref. Colangelo et al. (2022b), the calculation can be reduced to standard one-loop Passarino–Veltman functions $B_{0}$, $C_{0}$, $D_{0}$ Passarino and Veltman (1979) by inserting a dispersive representation of the pion form factor
@@
where the ${\mathcal{M}}_{k}$ represent a collection of loop functions and kinematic factors. In particular, using an unsubtracted dispersion relation (8) ensures that the result is manifestly UV finite. Moreover, we checked that IR singularities and the chiral logarithms agree with ChPT, by evaluating Eq. (Improved calculation of radiative corrections to $\boldsymbol{\tau\to\pi\pi\nu_{\tau}}$ decays) in the limit of a narrow resonance. Numerically, the evaluation becomes complicated by a superficial singularity at $s^{\prime}=s^{\prime\prime}$ in one of the IR-divergent $D_{0}$ functions Beenakker and Denner (1990), which disappears due to a factorization property at the border of the phase space—a cancellation that needs to be made manifest for a stable implementation. Based on the same factorization one can show explicitly that the resulting contribution to $G_{\text{EM}}(s)$ is finite at threshold, in agreement with the expectation from ChPT. More details of our dispersive result for $G_{\text{EM}}(s)$ are given in Ref. Colangelo et al. (2025).
@@
$f_{\text{loop}}^{\text{full}}(s,t)=f_{\text{loop}}^{\text{disp}}(s,t)-f_{\text{loop}}^{\text{disp}}(0,0)+f_{\text{loop}}^{\text{ChPT}}(0),$ (10)
@@
Finally, to produce numerical results we need to choose a value for the chiral LECs $X_{\ell}$ Descotes-Genon and Moussallam (2005), on which the final correction depends linearly, i.e., we can express our results at some value $\bar{X}_{\ell}(\mu)$ and restore the general result by means of Colangelo et al. (2025)
@@
Figure 3: Sample diagrams for the real corrections from the WZW anomaly as well as the exchange of vector and axial-vector meson resonances (indicated by the double lines). Otherwise notation as in Figs. 1 and 2.
Real corrections—For the real corrections we follow the general strategy from Refs. Cirigliano et al. (2001, 2002), including radiation off the $\tau$ and the charged pion, resonance contributions that saturate the LECs $L_{9}$, $L_{10}$ Gasser and Leutwyler (1984, 1985), and the prediction from the Wess–Zumino–Witten (WZW) anomaly Wess and Zumino (1971); Witten (1983), see Fig. 3. The leading corrections from Low’s theorem Low (1958) can be given in an analytic form, while the remainder has to be calculated numerically. In contrast to the leading Low contribution, which only diverges logarithmically at threshold, it behaves as $G_{\text{EM}}(s)\propto 1/(s-4M_{\pi}^{2})$, in such a way that a stable numerical evaluation becomes challenging close to threshold. This issue can be circumvented by a suitable change of integration variables Colangelo et al. (2025)—angular variables that yield stable expressions for amplitude times Jacobian—which allows us to evaluate both real and virtual contributions down to threshold in a robust manner.
@@
obtained from $\rho\to e^{+}e^{-}$, $\rho\to\pi\pi$, $K^{*}\to K\pi$ Colangelo et al. (2022d) and $a_{1}\to\pi\gamma$ Zielinski et al. (1984), respectively. The extraction of $F_{A}$ from $a_{1}\to\pi\gamma$ has been challenged in the literature mostly suggesting smaller values of $F_{A}$ Moussallam (1997); Knecht and Nyffeler (2001); Cirigliano et al. (2004); García-Martín and Moussallam (2010); Condo et al. (1993); Nozar et al. (2009), but the situation is far from conclusive. Hoping to shed some light on the issue, we attempted yet another indirect determination, by assuming that the $a_{1}\to\pi\gamma$ decay proceeds via $a_{1}\to\rho\pi\to\pi\gamma$ Colangelo et al. (2025); Hoferichter et al. (2017); Zanke et al. (2021), which gives values of $F_{A}=(0.07\ldots 0.13)\,\text{GeV}$, depending on the assumptions for the $a_{1}$ parameters. In view of the substantial uncertainty especially in $F_{A}$, we define our central values as the average of fits performed with either Eq. (Improved calculation of radiative corrections to $\boldsymbol{\tau\to\pi\pi\nu_{\tau}}$ decays) or Eq. (13) as input and assign the difference between the results as $1\sigma$ uncertainty. The resulting sizable error band should therefore encompass a reasonably broad set of vector-meson parameters, while including yet higher multiplets does not appear well motivated.
@@
where $\delta^{1}_{1}$ denotes the $P$-wave $\pi\pi$ phase shift, and calculate a first approximation for $G_{\text{EM}}(s)$. We then use the resulting correction in a fit of $f_{+}(s)$ to the $\tau$ data to obtain an improved $f_{+}(s)$. We recalculate $G_{\text{EM}}(s)$ with it and iterate the procedure until it converges after a few steps.
For the fit we use a form that combines the dispersive representation developed for $e^{+}e^{-}\to\pi^{+}\pi^{-}$ Colangelo et al. (2019); Colangelo et al. (2022c) with explicit resonance contributions for $\rho^{\prime}$, $\rho^{\prime\prime}$ Hoferichter et al. (2014, 2019); Hoferichter et al. (2023c, 2025d):
@@
Table 1: Summary of corrections to $a_{\mu}^{\text{HVP, LO}}[\pi\pi,\tau]$ in units of $10^{-10}$, see main text for the definition of the different rows. All results are given for the fit to Belle, Belle+ALEPH, and all four data sets. The first/second error refers to the experimental/theoretical uncertainty, the latter excluding the scheme ambiguity in the SD contribution and higher intermediate states.
@@
Ref. Davier et al. (2024) Ref. Castro et al. (2025) Ref. Aliberti et al. (2025a) This work
@@
Table 2: Comparison to previous work (notation as in Table 1). The SD uncertainty from Ref. Aliberti et al. (2025a) has been added to our results from Table 1, as has an additional uncertainty of $0.3$ units to capture higher intermediate states in the virtual contribution to $G_{\text{EM}}(s)$ Colangelo et al. (2025).
@@
changes by about three units (in $10^{-10}$). While the overall uncertainty is now dominated by the scheme dependence in the SD contribution, to be addressed in Ref. Cirigliano et al. (2025), the remaining uncertainty due to radiative corrections mainly reflects the role of higher intermediate states in the real and virtual contributions. Overall, the uncertainty due to $G_{\text{EM}}(s)$ has been reduced by almost a factor of three over the previously assigned error in Ref. Aliberti et al. (2025a), with a shift in central value by about $2.5\sigma$.
Besides improving the radiative corrections parameterized by $G_{\text{EM}}(s)$, our work also strongly motivates increased efforts in new measurements of the $\tau\to\pi\pi\nu_{\tau}$ spectral function, as possible at Belle II Altmannshofer et al. (2019). Indeed, our dispersive fits to the spectrum reveal that some tensions among the currently available data sets do exist, and at the same time we observe differences to previous evaluations, as visible in the context of $\tilde{a}_{\mu}$ in Table 2. Part of the difference might originate from the changes in $G_{\text{EM}}(s)$, but we also find that the constraints imposed by analyticity and unitarity result in a moderate tension between the low-energy part of the spectrum and the $\rho$ region, which tends to increase the integral for small values of $s$. An improved measurement of the $\tau\to\pi\pi\nu_{\tau}$ spectral function, profiting from the statistics available at Belle II, would therefore be extremely valuable.

### v2

@@ Abstract
and would allow for a test of the Standard Model at an unprecedented level of precision. Unfortunately, the limiting factor is represented by the precision of the current theory prediction Aliberti et al. (2025a); Aoyama et al. (2012); Hertzog and Hoferichter (2025); Volkov (2019, 2024); Aoyama et al. (2025); Parker et al. (2018); Morel et al. (2020); Fan et al. (2023); Czarnecki et al. (2003); Gnendiger et al. (2013); Lüdtke et al. (2025); Hoferichter et al. (2025a); Blum et al. (2018); Giusti et al. (2019); Borsányi et al. (2021); Lehner and Meyer (2020); Wang et al. (2023); Aubin et al. (2022); Cè et al. (2022); Alexandrou et al. (2023); Blum et al. (2023); Kuberski et al. (2024); Boccaletti et al. (2024); Spiegel and Lehner (2025); Blum et al. (2025a); Djukanovic et al. (2025); Alexandrou et al. (2025); Bazavov et al. (2025a); Bazavov et al. (2025b); Keshavarzi et al. (2020); Di Luzio et al. (2025); Kurz et al. (2014); Colangelo et al. (2015); Masjuan and Sánchez-Puertas (2017); Colangelo et al. (2017a, b); Hoferichter et al. (2018a, b); Eichmann et al. (2019); Bijnens et al. (2019); Leutgeb and Rebhan (2020); Cappiello et al. (2020); Masjuan et al. (2022); Bijnens et al. (2020, 2021); Danilkin et al. (2021); Stamen et al. (2022); Leutgeb et al. (2023); Hoferichter et al. (2023a); Hoferichter et al. (2024); Estrada et al. (2024); Deineka et al. (2025); Eichmann et al. (2025); Bijnens et al. (2025); Hoferichter et al. (2025b, c); Holz et al. (2025a, b); Cappiello et al. (2025); Colangelo et al. (2014); Blum et al. (2020); Chao et al. (2021, 2022); Blum et al. (2025b); Fodor et al. (2025)
@@
which is worse than the experimental one by about a factor of four. The main source of theory uncertainty is the leading-order (LO) hadronic-vacuum-polarization (HVP) contribution $a_{\mu}^{\text{HVP, LO}}$, which in Ref. Aliberti et al. (2025a) was estimated on the basis of lattice-QCD calculations, differing substantially from the previous $e^{+}e^{-}$-based result Aoyama et al. (2020); Davier et al. (2017); Keshavarzi et al. (2018); Colangelo et al. (2019); Hoferichter et al. (2019); Davier et al. (2020); Keshavarzi et al. (2020); Hoid et al. (2020). In fact, a data-driven estimate of this contribution was not even provided in Ref. Aliberti et al. (2025a) due to tensions in the $e^{+}e^{-}\to\text{hadrons}$ data base. The main rationale for this change was that after the CMD-3 measurement of the critical $e^{+}e^{-}\to\pi^{+}\pi^{-}$ channel Ignatov et al. (2024a, b) systematic tensions had increased to a level that could no longer be taken into account by a meaningful error inflation. Fortunately, there are a number of ongoing developments to rectify this situation Aliberti et al. (2025a); Colangelo et al. (2022a), including new data (such as the preliminary measurement by BaBar Polat (2026)) as well as improvements of radiative corrections and Monte-Carlo generators Campanario et al. (2019); Ignatov and Lee (2022); Colangelo et al. (2022b); Monnard (2021); Abbiendi et al. (2022); Lees et al. (2023); Budassi et al. (2025); Aliberti et al. (2025b); Fang et al. (2025); Budassi et al. (2026).
@@
A complementary avenue to shed light on the $2\pi$ contribution proceeds via $\tau\to\pi\pi\nu_{\tau}$ decays, in which case control over radiative corrections is arguably even more crucial. The $\tau$-based approach, pioneered in Ref. Alemany et al. (1998), relies on an isospin rotation from the $\pi^{\pm}\pi^{0}$ channel probed in $\tau$ decays back to $\pi^{+}\pi^{-}$ as required for the two-pion contribution $a_{\mu}^{\text{HVP, LO}}[\pi\pi]$. To make this work, first, the $\tau$-specific radiative corrections need to be removed from the measured spectral function and branching fraction, then isospin-breaking (IB) corrections need to be applied to the hadronic matrix elements, to account for the difference between the pion form factors $F_{\pi}^{V}(s)$ and $f_{+}(s)$ in the neutral and charged channel, respectively, and finally the $e^{+}e^{-}$-specific IB corrections need to be added back to match the conventional definition of $a_{\mu}^{\text{HVP, LO}}[\pi\pi]$. The latter can be extracted in a reliable way using a dispersive approach Colangelo et al. (2019, 2021); Colangelo et al. (2022c); Hoferichter et al. (2023b); Stoffer et al. (2023); Leplumey and Stoffer (2025) from the $e^{+}e^{-}\to\pi^{+}\pi^{-}$ data sets Achasov et al. (2006); Akhmetshin et al. (2007); Lees et al. (2012); Ablikim et al. (2016); Anastasi et al. (2018); Ignatov et al. (2024a), while the remaining IB corrections at present require model-dependent input Aliberti et al. (2025a) (see Ref. Colangelo et al. (2026a) for a first step towards a dispersive calculation), preventing the use of $\tau$ decays in the evaluation of $a_{\mu}^{\text{HVP, LO}}[\pi\pi]$.
@@
In this Letter, we address an improved calculation of the radiative corrections to $\tau\to\pi\pi\nu_{\tau}$, which presents a key challenge in the above program. So far, the evaluation of radiative corrections is based on chiral perturbation theory (ChPT) Cirigliano et al. (2001, 2002) in combination with resonance models Flores-Baéz et al. (2006, 2007); Davier et al. (2010); Miranda and Roig (2020); Davier et al. (2024); López Castro et al. (2025), which makes robust uncertainty quantification difficult. Most importantly, we consider structure-dependent virtual corrections based on a dispersive representation of the pion form factor, which were previously shown to be sizable in the case of $e^{+}e^{-}\to\pi^{+}\pi^{-}$ Ignatov and Lee (2022); Colangelo et al. (2022b). In this context, we pay particular attention to the matching to ChPT, to be able to make the connection to short-distance (SD) contributions as well as lattice QCD Bruno et al. (2018). We also revisit the calculation of real corrections, including the singular threshold behavior, and develop a method for a stable numerical evaluation of these corrections down to the two-pion threshold. Finally, we perform fits to the available $\tau$ spectral functions from Belle Fujikawa et al. (2008), ALEPH Schael et al. (2005); Davier et al. (2014), CLEO Anderson et al. (2000), and OPAL Ackerstaff et al. (1999), to assess the consequences for the IB corrections in a $\tau$-based evaluation of the $2\pi$ channel, $a_{\mu}^{\text{HVP, LO}}[\pi\pi,\tau]$.
@@
Figure 1: Virtual photonic corrections to $\tau\to\pi\pi\nu_{\tau}$ in ChPT. Solid lines denote leptons, dashed lines pions, and wiggly lines photons. Diagram $(a)$ refers to the chiral analog of the box diagram, while diagrams $(b)$ and $(c)$ originate from the five-particle vertex in ChPT. Not shown are self-energy diagrams and contact terms.
The SD contributions are subsumed into the electroweak (EW) correction factor $S_{\text{EW}}^{\pi\pi}=1+2\alpha/\pi\log(M_{Z}/m_{\tau})+\cdots=1.0233(3)(24)$ Sirlin (1982); Marciano and Sirlin (1986, 1988, 1993); Braaten and Li (1990); Erler (2004); Davier et al. (2003); Cirigliano et al. (2023), where the uncertainty reflects the $\mathcal{O}(\alpha/\pi)$ scheme ambiguity pointed out in Ref. Aliberti et al. (2025a). $f_{+}(s)$ denotes the pion form factor in the charged current, normalized as $f_{+}(0)=1$, and radiative corrections are expressed in terms of the electromagnetic (EM) correction factor
@@
which is the central object of this Letter. It is normalized in such a way that $G_{\text{EM}}(s)=1+\mathcal{O}(\alpha)$, and amounts to averaging the amplitudes $f_{\text{loop}}(s,t)$ and $g_{\text{rad}}(s,t)$ for virtual and real contributions, respectively, over the entire phase space. The variable $t$ represents the invariant mass squared of the $\tau$ and the charged pion, with integration boundaries $t_{\text{min}/\text{max}}(s)$, and the weight function reads
@@
This limitation can be overcome in a dispersive approach, in which case the diagrams in Fig. 1 need to be reinterpreted as unitarity diagrams. From this perspective, since diagrams $(b)$ and $(c)$ do not generate analytic singularities in the Mandelstam variables, only diagram $(a)$ needs to be considered further. Keeping in mind that the $\tau\nu_{\tau}\pi\pi^{0}$ vertex arises from integrating out the $W$ boson, this topology then takes the form shown in Fig. 2, and, accordingly, is referred to as “box diagram.” In general, its calculation requires knowledge of the full amplitude $\langle\pi\pi^{0}|j_{\text{EM}}^{\mu}j_{W}^{\nu}|0\rangle$ of EW and EM currents, which could be constructed dispersively in analogy to pion Compton scattering García-Martín and Moussallam (2010); Hoferichter et al. (2011); Moussallam (2013); Danilkin and Vanderhaeghen (2019); Hoferichter and Stoffer (2019); Danilkin et al. (2020). In practice, by far the most dominant correction will arise from the pion-pole contribution indicated in Fig. 2. Following the strategy of Ref. Colangelo et al. (2022b), the calculation can be reduced to standard one-loop Passarino–Veltman functions $B_{0}$, $C_{0}$, $D_{0}$ Passarino and Veltman (1979) by inserting a dispersive representation of the pion form factor
@@
where the ${\mathcal{M}}_{k}$ represent a collection of loop functions and kinematic factors. In particular, using an unsubtracted dispersion relation (8) ensures that the result is manifestly UV finite. Moreover, we checked that IR singularities and the chiral logarithms agree with ChPT, by evaluating Eq. (Improved calculation of radiative corrections to $\boldsymbol{\tau\to\pi\pi\nu_{\tau}}$ decays) in the limit of a narrow resonance. Numerically, the evaluation becomes complicated by a superficial singularity at $s^{\prime}=s^{\prime\prime}$ in one of the IR-divergent $D_{0}$ functions Beenakker and Denner (1990), which disappears due to a factorization property at the border of the phase space—a cancellation that needs to be made manifest for a stable implementation. Based on the same factorization one can show explicitly that the resulting contribution to $G_{\text{EM}}(s)$ is finite at threshold, in agreement with the expectation from ChPT. More details of our dispersive result for $G_{\text{EM}}(s)$ are given in Ref. Colangelo et al. (2026b), which contains all the details concerning the analyses presented in this Letter.
@@
$f_{\text{loop}}^{\text{full}}(s,t)=f_{\text{loop}}^{\text{disp}}(s,t)-f_{\text{loop}}^{\text{disp}}(0,0)+f_{\text{loop}}^{\text{ChPT}}(0,0),$ (10)
@@
Finally, to produce numerical results we need to choose a value for the chiral LECs $X_{\ell}$ Descotes-Genon and Moussallam (2005), on which the final correction depends linearly, i.e., we can express our results at some value $\bar{X}_{\ell}(\mu)$ and restore the general result by means of Colangelo et al. (2026b)
@@
Figure 3: Sample diagrams for the real corrections from the WZW anomaly [diagram $(a)$] as well as the exchange of vector [diagram $(b)$] and axial-vector [diagram $(c)$] meson resonances (indicated by the double lines). Otherwise notation as in Figs. 1 and 2.
Real corrections—For the real corrections we follow the general strategy from Refs. Cirigliano et al. (2001, 2002), including radiation off the $\tau$ and the charged pion, resonance contributions that saturate the LECs $L_{9}$, $L_{10}$ Gasser and Leutwyler (1984, 1985), and the prediction from the Wess–Zumino–Witten (WZW) anomaly Wess and Zumino (1971); Witten (1983), see Fig. 3. The leading corrections from Low’s theorem Low (1958) can be given in an analytic form, while the remainder has to be calculated numerically. In contrast to the leading Low contribution, which only diverges logarithmically at threshold, it behaves as $G_{\text{EM}}(s)\propto 1/(s-4M_{\pi}^{2})$, in such a way that a stable numerical evaluation becomes challenging close to threshold. This issue can be circumvented by a suitable change of integration variables Colangelo et al. (2026b)—angular variables that yield stable expressions for amplitude times Jacobian—which allows us to evaluate both real and virtual contributions down to threshold in a robust manner.
@@
obtained from $\rho\to e^{+}e^{-}$, $\rho\to\pi\pi$, $K^{*}\to K\pi$ Colangelo et al. (2022d), and $a_{1}\to\pi\gamma$ Zielinski et al. (1984), respectively. The extraction of $F_{A}$ from $a_{1}\to\pi\gamma$ has been challenged in the literature mostly suggesting smaller values of $F_{A}$ Moussallam (1997); Knecht and Nyffeler (2001); Cirigliano et al. (2004); García-Martín and Moussallam (2010); Condo et al. (1993); Nozar et al. (2009), but the situation is far from conclusive. Hoping to shed some light on the issue, we attempted yet another indirect determination, by assuming that the $a_{1}\to\pi\gamma$ decay proceeds via $a_{1}\to\rho\pi\to\pi\gamma$ Colangelo et al. (2026b); Hoferichter et al. (2017); Zanke et al. (2021), which gives values of $F_{A}=(0.07\ldots 0.13)\,\text{GeV}$, depending on the assumptions for the $a_{1}$ parameters. In view of the substantial uncertainty especially in $F_{A}$, we define our central values as the average of fits performed with either Eq. (Improved calculation of radiative corrections to $\boldsymbol{\tau\to\pi\pi\nu_{\tau}}$ decays) or Eq. (13) as input and assign the difference between the results as $1\sigma$ uncertainty. The resulting sizable error band should therefore encompass a reasonably broad set of vector-meson parameters, while including yet higher multiplets does not appear well motivated.
@@
where $\delta^{1}_{1}$ denotes the $P$-wave $\pi\pi$ phase shift, and calculate a first approximation for $G_{\text{EM}}(s)$. We then use the resulting correction in a fit of $f_{+}(s)$ to the $\tau$ data, see End Matter for more details of the fit function (including a conformal polynomial with $N$ parameters to describe inelastic effects as well as explicit parameterizations of the $\rho^{\prime}$, $\rho^{\prime\prime}$ resonances). We recalculate $G_{\text{EM}}(s)$ with the improved $f_{+}(s)$ obtained from the fit and iterate the procedure until it converges after a few steps.
@@
Table 1: Summary of corrections to $a_{\mu}^{\text{HVP, LO}}[\pi\pi,\tau]$ in units of $10^{-10}$, see main text for the definition of the different rows. All results are given for the fit to Belle, Belle+ALEPH, and all four data sets. The first (second) error refers to the experimental (theoretical) uncertainty, the latter excluding the scheme ambiguity in the SD contribution and higher intermediate states.
@@
Ref. Davier et al. (2024) Ref. López Castro et al. (2025) Ref. Aliberti et al. (2025a) This Letter
@@
Table 2: Comparison to previous work (notation as in Table 1). The SD uncertainty from Ref. Aliberti et al. (2025a) has been added to our results from Table 1, as has an additional uncertainty of $0.3$ units to capture higher intermediate states in the virtual contribution to $G_{\text{EM}}(s)$ Colangelo et al. (2026b).
@@
changes by about three units (in $10^{-10}$). While the overall uncertainty is now dominated by the scheme dependence in the SD contribution, addressed in Ref. Cirigliano et al. (2026), the remaining uncertainty due to radiative corrections mainly reflects the role of higher intermediate states in the real and virtual contributions. Overall, the uncertainty due to $G_{\text{EM}}(s)$ has been reduced by almost a factor of three over the previously assigned error in Ref. Aliberti et al. (2025a), with a shift in central value by about $2.5\sigma$.
Besides improving the radiative corrections parameterized by $G_{\text{EM}}(s)$, our work also strongly motivates increased efforts in new measurements of the $\tau\to\pi\pi\nu_{\tau}$ spectral function, as possible at Belle II Altmannshofer et al. (2019). Indeed, our dispersive fits to the spectrum reveal that some tensions among the currently available data sets do exist, and at the same time we observe differences to previous evaluations, as visible in the context of $\tilde{a}_{\mu}$ in Table 2. Part of the difference might originate from the changes in $G_{\text{EM}}(s)$ dat , but we also find that the constraints imposed by analyticity and unitarity result in a moderate tension between the low-energy part of the spectrum and the $\rho$ region, which tends to increase the integral for small values of $s$. An improved measurement of the $\tau\to\pi\pi\nu_{\tau}$ spectral function, profiting from the statistics available at Belle II, would therefore be extremely valuable.